The sum of two number is 98. If three times the smaller number is subtracted from the larger number, the result is 10. Find the two numbers.
Answer by macston(5194)

.
S=smaller number; L=larger number
L+S=98
L=98-S Use this to substitute for L
L-3S=10 Substitute for L from above.
98-S-3S=10 Subtract 98 from each side.
-4S=-88 Divide each side by -4
S=22 ANSWER 1: The smaller number is 22.
L=98-S=98-22=76 ANSWER 2: The larger number is 76.
.
CHECK:
L-3S=10
76-3(22)=10
76-66=10
10=10